OpenStudy (saylilbaby):

simplify: radical sign 10a*radical sign 30a

OpenStudy (saylilbaby):

@StudyGurl14 @austin11111 @texaschic101

OpenStudy (epoweritheta):

u mean \[\sqrt{10a} \times \sqrt{30a}\]

OpenStudy (saylilbaby):

yes @epoweritheta

OpenStudy (texaschic101):

10a radical sign 3

OpenStudy (saylilbaby):

how do i do a radical sign @epoweritheta

OpenStudy (epoweritheta):

 now multiply under the radical